# Fusion Systems related to polynomial representations of $\operatorname{SL}_{2}(q)$

### v1

@@ 2. The structure of polynomial $p$-groups and their automorphisms
Let $q=p^{m}$ and $S\in\{S_{n}(q),S_{\Lambda}(q)\}$. Then for $B\in\mathcal{B}(S)$ and $C\in\mathcal{C}(S)$, the following hold:
@@
    if $B,B^{\prime}\in\mathcal{B}(S)$ are such that $B\neq B^{\prime}$ then $B\cap B^{\prime}=Z(S)$; and
@@
    if $C,C^{\prime}\in\mathcal{C}(S)$ are such that $C\neq C^{\prime}$ then $C\cap C^{\prime}=Z_{2}(S)$.
@@
Let $B,B^{\prime},C,C^{\prime}$ be as in the statement of the lemma. By Lemmas 2.4 and 2.12, we have that $B=C_{S}(b)$ for $b\in B\setminus Z(S)$ and $|C_{C}(c)|=|C_{S}(c)|=q^{2}d$ for $c\in C\setminus Z_{2}(S)$. Since $Z(S)\leq B\cap C$, if there is $b\in(B\cap C)\setminus Z(S)$, then $b\in C\setminus Z_{2}(S)$ and so $B\leq C_{S}(b)=C_{C}(b)$. Comparing orders we obtain that $B=C_{C}(b)$, proving (1). Next note that any $b\in B\cap B^{\prime}$ is centralised by both $B$ and $B^{\prime}$ and so if $B\neq B^{\prime}$ then necessarily $B\cap B^{\prime}=Z(S)$. This proves (2). Finally to prove (3) note that $Z_{2}(S)\leq C\cap C^{\prime}$ and if $c\in(C\cap C^{\prime})\setminus Z_{2}(S)$, we must have $C=C_{S}(c)Z_{2}(S)=C_{C^{\prime}}(c)Z_{2}(S)=C^{\prime}$, a contradiction. ∎
@@
To close this section, we focus our attention on the groups $S=S_{n}(q)$ for $11\leqslant n\leqslant p-$ 1. For $X\unlhd S$, define
@@
Suppose $\phi\in\mathrm{Aut}(S)$ has $p^{\prime}$-order and $(\phi)\delta=(1,1)$. Then $\phi$ centralises $S/V$ and $Z(S)=C_{V}(S)$. The Thompson A×$\times$×B-Lemma [Asc86, (24.2)] now implies that $\phi$ centralises $V$ and coprime action yields $\phi=1$. Hence $\ker\delta$ is a $p$-group. Since $\mathrm{Inn}(S)$ centralises $S/V$ and $Z(S)$, it follows that $\mathrm{Inn}(S)\leq\ker(\delta)$. ∎
Notice that $\mathrm{Aut}_{B}(S)\delta=\mathrm{Aut}_{\Sigma}(S)\delta$ has order $(q-1)^{2}.m_{p^{\prime}}$, where $m_{p^{\prime}}$ denotes the $p^{\prime}$-part of $m$. Identifying $S/V$ and $Z(S)$ with $\mathbb{K}$, taking $t=(\phi,\lambda,\begin{pmatrix}\mu&0\\ 0&\nu\end{pmatrix})\in\Sigma$ and letting $c_{t}\in\mathrm{Aut}(S)$ be the automorphism of $S$ induced by conjugation by $t$, we obtain
@@ 3. A new family of exotic fusion systems
Adopting the methodology in [CP10], the objective of this section is to describe exotic fusion systems supported on $S_{\Lambda}(q)$ for various $q$. Fix $\mathbb{K}$ of size $q=p^{m}$ for $p$ an odd prime. Set $V=\Lambda(q)$ and $S=S_{\Lambda}(q)$.
Recall from Section 2 that we write $P^{*}=D^{*}\ltimes V$ and $P=D\ltimes V$ so that $P\leq P^{*}$. Then elements of $P^{*}$ are regarded as tuples and the normaliser of a Sylow $p$-subgroup $S$ of $P^{*}$ is
@@
Define
$R=\left\langle Z(S),(1,1,\left(\begin{smallmatrix}1&0\\ \lambda&1\end{smallmatrix}\right),0)\mid\lambda\in\mathbb{K}\right\rangle$
and observe that $R\in\mathcal{B}(S)$ and
$N_{S}(R)=R\langle\overline{x^{2}y^{p-2}}\rangle_{\mathbb{K}}=RZ_{2}(S).$
Hence, $R$ is an elementary abelian $p$-group of order $q^{3}$ and $|N_{S}(R)|=q^{4}$ by Lemma 2.12(2).
@@
For a proof, see Proposition A.1 in the appendix. ∎
@@
Thus $F^{*}$ and $F$ are infinite groups with subgroups isomorphic to $P^{*}$ and $P^{*}_{R}$, and $P$ and $P_{R}$ respectively, which intersect in a subgroup isomorphic to $N_{P^{*}}(R)$, $N_{P}(R)$ respectively. We identify $P^{*},P_{R}^{*}$ and $S$ as subgroups of $F^{*}$, and then define
$\mathcal{F}_{\Lambda}^{*}(q)=\mathcal{F}_{S}(F^{*})$
to be the fusion system of $F^{*}$ on $S$. Similarly, we identify $P,P_{R}$ and $S$ as subgroups of $F$, and then define
$\mathcal{F}_{\Lambda}(q)=\mathcal{F}_{S}(F)$
to be the fusion system of $F$ on $S$. Clearly, $\mathcal{F}_{\Lambda}(q)$ is a fusion subsystem of $\mathcal{F}_{\Lambda}^{*}(q)$. Our main result of this section is:
@@
We prove Proposition 3.2 only for $\mathcal{F}_{\Lambda}^{*}(q)$ but the proof for $\mathcal{F}_{\Lambda}(q)$ is the same. We intend to apply [PS18, Theorem 2.5] with $G=F^{*}$. For convenience we set $M^{*}=P^{*}_{R}$. Let $\Gamma$ be the coset graph of the amalgam $F$. Thus $\Gamma$ has vertex set the right cosets of $P^{*}$ and $M^{*}$ in $F^{*}$ and two vertices are incident if and only if their intersection is non-empty and they are not equal. We know that $F^{*}$ acts edge transitively by right multiplication on $\Gamma$. Since $F^{*}$ is the free amalgamated product of $P^{*}$ and $M^{*}$, $\Gamma$ is a tree. For more details see [PR02, Section 4.1].
@@
We now use [AKO11, Theorem I.7.7] to show that $O^{p^{\prime}}(\mathcal{F}_{\Lambda}^{*}(q))=\mathcal{F}_{\Lambda}(q)$, adopting the notation from [AKO11]. We have that $\mathcal{E}_{0}$ contains $O^{p^{\prime}}(\mathrm{Aut}_{P^{*}}(V))=O^{p^{\prime}}(\mathrm{Aut}_{P}(V))$ which contains the conjugation map which corresponds to $(1,1,\left(\begin{smallmatrix}a&0\\ 0&a^{-1}\end{smallmatrix}\right),0)\in P^{*}$, where $a\in\mathbb{K}^{*}$. Moreover, $\mathcal{E}_{0}$ also contains the conjugation map determined by $d=\mathrm{diag}(1,\zeta^{-1},\zeta,1)\in P_{R}$ for all $\zeta\in\mathbb{K}^{*}$ on $R$. By Proposition 3.1 such maps correspond to conjugations by elements from $C_{D^{*}}(V)\left\langle\left(1,\zeta^{p},\left(\begin{matrix}1&0\\ 0&\zeta^{-1}\end{matrix}\right),0\right)\mid\zeta\in\mathbb{K}^{*}\right\rangle$ where we recognise $C_{D^{*}}(V)=\left\{\left(1,a^{-p},\begin{pmatrix}a&0\\ 0&a\end{pmatrix},0\right)\mid a\in\mathbb{K}^{*}\right\}$ as a subgroup of $P^{*}$. Indeed, the conjugation map determined to $d$ lifts to a faithful automorphism of $S$ and generates a subgroup of order $(q-1)$ in $\mathrm{Aut}_{\mathcal{F}}(S)$. Since this subgroup intersects $\langle(1,1,\left(\begin{smallmatrix}a&0\\ 0&a^{-1}\end{smallmatrix}\right),0)\rangle$ trivially, we deduce that $\mathrm{Aut}_{\mathcal{F}_{\Lambda}^{*}(q)}^{0}(S)=\mathrm{Aut}_{\mathcal{F}_{\Lambda}(q)}^{0}(S)=\mathrm{Aut}_{\mathcal{F}_{\Lambda}(q)}(S)$ and now $\mathcal{F}_{\Lambda}(q)=O^{p^{\prime}}(\mathcal{F}_{\Lambda}^{*}(q))$ by [AKO11, Theorem I.7.7 (d)]. ∎
@@
Set $\mathcal{F}=\mathcal{F}_{\Lambda}(q)$ so that $\mathcal{F}=O^{p^{\prime}}(\mathcal{F})$. Suppose that $X$ is strongly closed in $S$. Then $X$ is normal in $S$ and $C_{V}(S)\cap X\neq 1$. Let $z\in C_{V}(S)\cap X$ be non-trivial. Then $\langle z^{\mathrm{Aut}_{\mathcal{F}}(V)}\rangle$ contains the invariant subgroup $W$ of $V$ of order $q^{p-1}$. Thus we may assume that $z$ is in $C_{W}(S)=\langle\overline{xy^{p-1}}\rangle_{\mathbb{K}}$ Lemma 2.12(1). But then $X\geq\langle z^{\mathrm{Aut}_{\mathcal{F}}(R)}\rangle\not\leq V$. Let $w\in X\setminus V$. Then $X\geq[V,X]\geq[V,w]$. As $|V/[V,w]|=q^{2}$ by Lemma 2.12(3), and since $[V,w]\not\leq W$, we have $\langle[V,X]^{\mathrm{Aut}_{\mathcal{F}}(V)}\rangle=V$. Thus $V<X$. Finally, $X=\langle X^{\mathrm{Aut}_{\mathcal{F}}(S)}\rangle=S$. Then [AKO11, Theorem II.9.8(d)] yields that $\mathcal{F}$ is simple. ∎
@@
If $F^{*}(G)$ is an alternating group $\mathrm{Alt}(n)$, then, as $S$ is not a direct product, $n=p^{k}$. But in this case the Sylow $p$-subgroups of $\mathrm{Alt}(n)$ are iterated wreath product which have cyclic centre.
@@
We have shown that $F^{*}(G)$ is a Lie type group defined in characteristic $r$, $r\neq p$. Suppose that $G$ is not an exceptional group with $p\in\{3,5\}$. Then [GLS98, Theorem 4.10.3(e)] first yields that $V$ is contained in a maximal rank abelian subgroup of $S$, and then that $R$ is conjugate to a subgroup of $V$. This contradicts the fact that $R$ is $\mathcal{F}$-centric. Hence, it remains to consider the exceptional groups when $p\in\{3,5\}$. If $|S/V|=p$ then we appeal to [COS17, Theorem 2.8] for a contradiction.

### v2

@@ 2. The structure of polynomial $p$-groups and their automorphism group
Let $q=p^{m}$ and $S\in\{S_{n}(q),S_{\Lambda}(q)\}$. Then for $B_{0}\in\mathcal{B}(S)$ and $C_{0}\in\mathcal{C}(S)$, the following hold:
@@
    if $B_{0},B_{1}\in\mathcal{B}(S)$ are such that $B_{0}\neq B_{1}$ then $B_{0}\cap B_{1}=Z(S)$; and
@@
    if $C_{0},C_{1}\in\mathcal{C}(S)$ are such that $C_{0}\neq C_{1}$ then $C_{0}\cap C_{1}=Z_{2}(S)$.
@@
Let $B_{0},B_{1},C_{0},C_{1}$ be as in the statement of the lemma. By Lemmas 2.8 and 2.16, we have that $B_{0}=C_{S}(b)$ for $b\in B_{0}\setminus Z(S)$ and $|C_{C_{0}}(c)|=|C_{S}(c)|=q^{2}d$ for $c\in C_{0}\setminus Z_{2}(S)$. Since $Z(S)\leq B_{0}\cap C_{0}$, if there is $b\in(B_{0}\cap C_{0})\setminus Z(S)$, then $b\in C_{0}\setminus Z_{2}(S)$ and so $B_{0}\leq C_{S}(b)=C_{C_{0}}(b)$. Comparing orders we obtain that $B_{0}=C_{C_{0}}(b)$, proving (1). Next note that any $b\in B_{0}\cap B_{1}$ is centralised by both $B_{0}$ and $B_{1}$ and so if $B_{0}\neq B_{1}$ then necessarily $B_{0}\cap B_{1}=Z(S)$. This proves (2). Finally to prove (3) note that $Z_{2}(S)\leq C_{0}\cap C_{1}$ and if $c\in(C_{0}\cap C_{1})\setminus Z_{2}(S)$, we must have $C_{0}=C_{S}(c)Z_{2}(S)=C_{C_{1}}(c)Z_{2}(S)=C_{1}$, a contradiction. ∎
@@
To close this section, we focus our attention on the groups $S=S_{n}(q)$ for $1\leqslant n\leqslant p-1$.
Notation 2.25.
For $X\unlhd S$, define
@@
[…]
Notice that $\mathrm{Aut}_{B}(S)\delta=\mathrm{Aut}_{\Sigma}(S)\delta$ has order $(q-1)^{2}.m_{p^{\prime}}$, where $m_{p^{\prime}}$ denotes the $p^{\prime}$-part of $m$. Recall the group $\Sigma$ from 2.2. Identifying $S/V$ and $Z(S)$ with $\mathbb{K}$, taking $t=(\phi,\lambda,\begin{pmatrix}\mu&0\\ 0&\nu\end{pmatrix})\in\Sigma$ and letting $c_{t}\in\mathrm{Aut}(S)$ be the automorphism of $S$ induced by conjugation by $t$, we obtain
@@ 3. A new family of exotic fusion systems
Adopting the methodology in [CP10], the objective of this section is to describe exotic fusion systems supported on $S_{\Lambda}(q)$ for various $q$. To prove these examples are saturated, we will apply the following result from [PS18]:
Theorem 3.1.
Let $\mathcal{A}=(G_{1}\geq G_{12}\leq G_{2})$ be an amalgam of finite groups, assume that $\operatorname{Syl}_{p}(G_{12})\subseteq\operatorname{Syl}_{p}(G_{2})$ and fix $S_{i}\in\operatorname{Syl}_{p}(G_{i})$ with $S_{2}\leq S_{1}$. Assume that $G=G_{1}*_{G_{12}}G_{2}$ is the universal completion of $\mathcal{A}$ and write $\Gamma=\Gamma(G,G_{1},G_{2},G_{12})$ for the coset graph. Suppose that:
  1. (a)
    for all $\mathcal{F}_{S_{1}}(G)$-centric subgroups $P$ of $S_{1}$, $\Gamma^{P}$ is finite; and
  2. (b)
    each $\mathcal{F}_{S_{i}}(G_{i})$-essential subgroup is $\mathcal{F}_{S_{1}}(G)$-centric.
Then $\mathcal{F}_{S_{1}}(G)$ is saturated.
Proof.
See [PS18, Theorem 2.5]. ∎
Let $\Gamma$ be the coset graph of the amalgam $\mathcal{A}$ in Theorem 3.1. Thus $\Gamma$ has vertex set the right cosets of $G_{1}$ and $G_{2}$ in $G$ and two vertices are incident if and only if their intersection is non-empty and they are not equal. We know that $G$ acts edge transitively by right multiplication on $\Gamma$. Since $G$ is the free amalgamated product of $G_{1}$ and $G_{2}$, $\Gamma$ is a tree. For more details see [PR02, Section 4.1].
Fix $\mathbb{K}$ of size $q=p^{m}$ for $p$ an odd prime. Set $V=\Lambda(q)$ and $S=S_{\Lambda}(q)$. Recall from Notation 2.5 that we write $P_{\Lambda}^{*}(q)=D^{*}\ltimes V$ and $P_{\Lambda}(q)=D\ltimes V$ so that $P_{\Lambda}(q)\leq P_{\Lambda}^{*}(q)$.
For this section, write $P^{*}:=P_{\Lambda}^{*}(q)$ and $P:=P_{\Lambda}(q)$. Then elements of $P^{*}$ are regarded as tuples and the normaliser of a Sylow $p$-subgroup $S$ of $P^{*}$ is
@@
Recall from Notations 2.2 and 2.21 that $R=UZ(S)=\left\langle Z(S),(1,1,\left(\begin{smallmatrix}1&0\\ \lambda&1\end{smallmatrix}\right),0)\mid\lambda\in\mathbb{K}\right\rangle\in\mathcal{B}(S)$. Then $N_{S}(R)=R\langle\overline{x^{2}y^{p-2}}\rangle_{\mathbb{K}}=RZ_{2}(S)$, $R$ is an elementary abelian $p$-group of order $q^{3}$ and $|N_{S}(R)|=q^{4}$ by Lemma 2.16(2).
@@
$\left(\phi\Phi,\theta^{p^{l}}\Pi,\left(\begin{matrix}a^{p^{l}}&0\\ c^{p^{l}}&b^{p^{l}}\end{matrix}\right)\left(\begin{matrix}\alpha&0\\ \kappa&\beta\end{matrix}\right),\left(\lambda^{p^{l}}\overline{y^{p}}+\mu^{p^{l}}\overline{xy^{p-1}}+\nu^{p^{l}}\overline{x^{2}y^{p-2}}\right)\cdot\left(\Pi,\left(\begin{matrix}\alpha&0\\ \kappa&\beta\end{matrix}\right)\right)+\chi\overline{y^{p}}+\tau\overline{xy^{p-1}}+\eta\overline{x^{2}y^{p-2}}\right)$
$\left(\phi\Phi,\theta^{p^{l}}\Pi,\left(\begin{matrix}a^{p^{l}}\alpha&0\\ c^{p^{l}}\alpha+\kappa b^{p^{l}}&b^{p^{l}}\beta\end{matrix}\right),(\lambda^{p^{l}}\Pi^{-1}\beta^{-p}+\chi)\overline{y^{p}}+(\nu^{p^{l}}\Pi^{-1}\kappa\alpha^{-2}\beta^{1-p}+\mu^{p^{l}}\Pi^{-1}\alpha^{-1}\beta^{1-p}+\tau)\overline{xy^{p-1}}+(\nu^{p^{l}}\Pi^{-1}\alpha^{-2}\beta^{2-p}+\eta)\overline{x^{2}y^{p-2}}\right).$
On the other hand,
$\left(\phi\Phi,\left(\begin{smallmatrix}\theta(ab^{p})^{p^{l}}&0&0&(\theta ab^{p}\lambda)^{p^{l}}\\ 0&(\theta a^{2}b^{p-1})^{p^{l}}&(\theta a^{2}b^{p-1}\nu)^{p^{l}}&(\theta a^{2}b^{p-1}\mu)^{p^{l}}\\ 0&0&b^{p^{l}}&c^{p^{l}}\\ 0&0&0&a^{p^{l}}\end{smallmatrix}\right)\left(\begin{smallmatrix}\Pi\alpha\beta^{p}&0&0&\Pi\alpha\beta^{p}\chi\\ 0&\Pi\alpha^{2}\beta^{p-1}&\Pi\alpha^{2}\beta^{p-1}\eta&\Pi\alpha^{2}\beta^{p-1}\tau\\ 0&0&\beta&\kappa\\ 0&0&0&\alpha\end{smallmatrix}\right)\right)$
$\left(\phi\Phi,\left(\begin{smallmatrix}(\theta ab^{p})^{p^{l}}\Pi\alpha\beta^{p}&0&0&(\theta ab^{p})^{p^{l}}\Pi\alpha\beta^{p}\chi+(\theta ab^{p}\lambda)^{p^{l}}\alpha\\ 0&(\theta a^{2}b^{p-1})^{p^{l}}\Pi\alpha^{2}\beta^{p-1}&(\theta a^{2}b^{p-1})^{p^{l}}\Pi\alpha^{2}\beta^{p-1}\eta+(\theta a^{2}b^{p-1}\nu)^{p^{l}}\beta&(\theta a^{2}b^{p-1})^{p^{l}}\Pi\alpha^{2}\beta^{p-1}\tau+(\theta a^{2}b^{p-1}\nu)^{p^{l}}\kappa+(\theta a^{2}b^{p-1}\mu)^{p^{l}}\alpha\\ 0&0&b^{p^{l}}\beta&b^{p^{l}}\kappa+\alpha c^{p^{l}}\\ 0&0&0&a^{p^{l}}\alpha\end{smallmatrix}\right)\right).$
It is elementary to check that (1) and (2) hold. ∎
@@
Thus $F^{*}$ and $F$ are infinite groups with subgroups isomorphic to $P^{*}$ and $P^{*}_{R}$, and $P$ and $P_{R}$ respectively, which intersect in a subgroup isomorphic to $N_{P^{*}}(R)$, $N_{P}(R)$ respectively. We identify $P^{*},P_{R}^{*}$ and $P,P_{R}$ as subgroups of $F^{*}$ and $F$ respectively. In particular $S$ is identified as a subgroup of both $F^{*}$ and $F$.
Notation 3.3.
Define the fusion systems on $S=S_{\Lambda}(q)$:
[…]
Clearly, $\mathcal{F}_{\Lambda}(q)$ is a fusion subsystem of $\mathcal{F}_{\Lambda}^{*}(q)$. Our main result of this section is:
@@
We prove Proposition 3.4 only for $\mathcal{F}_{\Lambda}^{*}(q)$ but the proof for $\mathcal{F}_{\Lambda}(q)$ is the same. We intend to apply Theorem 3.1 with $G=F^{*}$. For convenience we set $M^{*}=P^{*}_{R}$. Thus, as above, $\Gamma$ has vertex set the right cosets of $P^{*}$ and $M^{*}$ in $F^{*}$ and two vertices are incident if and only if their intersection is non-empty and they are not equal. Then $F^{*}$ acts edge transitively by right multiplication on $\Gamma$, and $\Gamma$ is a tree.
@@
We now use [AKO11, Theorem I.7.7] to show that $O^{p^{\prime}}(\mathcal{F}_{\Lambda}^{*}(q))=\mathcal{F}_{\Lambda}(q)$, adopting the notation from [AKO11]. We have that $\mathcal{E}_{0}$ contains $O^{p^{\prime}}(\mathrm{Aut}_{P^{*}}(V))=O^{p^{\prime}}(\mathrm{Aut}_{P}(V))$ which contains the conjugation map which corresponds to $(1,1,\left(\begin{smallmatrix}a&0\\ 0&a^{-1}\end{smallmatrix}\right),0)\in P^{*}$, where $a\in\mathbb{K}^{*}$. Moreover, $\mathcal{E}_{0}$ also contains the conjugation map determined by $d=\mathrm{diag}(1,\zeta^{-1},\zeta,1)\in P_{R}$ for all $\zeta\in\mathbb{K}^{*}$ on $R$. By Proposition 3.2 such maps correspond to conjugations by elements from $C_{D^{*}}(V)\left\langle\left(1,\zeta^{p},\left(\begin{matrix}1&0\\ 0&\zeta^{-1}\end{matrix}\right),0\right)\mid\zeta\in\mathbb{K}^{*}\right\rangle$ where we recognise $C_{D^{*}}(V)=\left\{\left(1,a^{-p},\begin{pmatrix}a&0\\ 0&a\end{pmatrix},0\right)\mid a\in\mathbb{K}^{*}\right\}$ as a subgroup of $P^{*}$. Indeed, the conjugation map determined to $d$ lifts to a faithful automorphism of $S$ and generates a subgroup of order (q1)(q-1) in $\mathrm{Aut}_{\mathcal{F}}(S)$. Since this subgroup intersects $\langle(1,1,\left(\begin{smallmatrix}a&0\\ 0&a^{-1}\end{smallmatrix}\right),0)\rangle$ trivially, we deduce that $\mathrm{Aut}_{\mathcal{F}_{\Lambda}^{*}(q)}^{0}(S)=\mathrm{Aut}_{\mathcal{F}_{\Lambda}(q)}^{0}(S)=\mathrm{Aut}_{\mathcal{F}_{\Lambda}(q)}(S)$ and now $\mathcal{F}_{\Lambda}(q)=O^{p^{\prime}}(\mathcal{F}_{\Lambda}^{*}(q))$ by [AKO11, Theorem I.7.7 (d)]. ∎
@@
Set $\mathcal{F}=\mathcal{F}_{\Lambda}(q)$ so that $\mathcal{F}=O^{p^{\prime}}(\mathcal{F})$. Suppose that $X$ is strongly closed in $S$. Then $X$ is normal in $S$ and $C_{V}(S)\cap X\neq 1$. Let $z\in C_{V}(S)\cap X$ be non-trivial. Then $\langle z^{\mathrm{Aut}_{\mathcal{F}}(V)}\rangle$ contains the invariant subgroup $W$ of $V$ of order $q^{p-1}$. Thus we may assume that $z$ is in $C_{W}(S)=\langle\overline{xy^{p-1}}\rangle_{\mathbb{K}}$ Lemma 2.16(1). But then $X\geq\langle z^{\mathrm{Aut}_{\mathcal{F}}(R)}\rangle\not\leq V$. Let $w\in X\setminus V$. Then $X\geq[V,X]\geq[V,w]$. As $|V/[V,w]|=q^{2}$ by Lemma 2.16(3), and since $V,w]\not\leq W$, we have $\langle[V,X]^{\mathrm{Aut}_{\mathcal{F}}(V)}\rangle=V$. Thus $V<X$. Finally, $X=\langle X^{\mathrm{Aut}_{\mathcal{F}}(S)}\rangle=S$. Then [AKO11, Theorem II.9.8(d)] yields that $\mathcal{F}$ is simple. ∎
@@
If $F^{*}(G)$ is an alternating group $\mathrm{Alt}(n)$, then, as $S$ is not a direct product, we see that $n=p^{k}$. But in this case the Sylow $p$-subgroups of $\mathrm{Alt}(n)$ are iterated wreath products which have cyclic centre.
@@
We have shown that $F^{*}(G)$ is a Lie type group defined in characteristic $r$, where $r\neq p$. Suppose that $G$ is not an exceptional group with $p\in\{3,5\}$. Then [GLS98, Theorem 4.10.3(e)] first yields that $V$ is contained in a maximal rank abelian subgroup of $S$, and then that $R$ is conjugate to a subgroup of $V$. This contradicts the fact that $R$ is $\mathcal{F}$-centric. Hence, it remains to consider the exceptional groups when $p\in\{3,5\}$. If $|S/V|=p$ then we appeal to [COS17, Theorem 2.8] for a contradiction.
@@
Combining the results in this section, we have demonstrated the following theorem.
Theorem 3.9.
For $p$ an odd and $q=p^{m}$, the group $S_{\Lambda}(q)$ supports a saturated fusion system $\mathcal{F}^{*}=\mathcal{F}_{\Lambda}^{*}(q)$. Moreover, $O^{p^{\prime}}(\mathcal{F}^{*})$ is simple and all $p^{\prime}$-index subsystems of $\mathcal{F}^{*}$ are exotic.
Remark 3.10.
Assume that $p$ is odd, $q=p^{m}$, and let $\mathcal{F}$ be a fusion subsystem of $p^{\prime}$-index in $\mathcal{F}^{*}_{\Lambda}(q)$ on $S=S_{\Lambda}(q)$.
When $q=p$, the group $S$ has an elementary abelian subgroup of index $p$, and the fusion systems in Theorem 3.9 occur among the examples considered in [COS17] (see the second and seventh rows of [COS17, Table 4.1] and [COS17, Proposition 4.1]). On the other hand, when $q>p$ these examples appear to be new.
At the time of writing, apart from a handful of further examples mentioned below, the majority of known exotic fusion systems are described in [GP25]. Broadly speaking, these fall into six families, which we now consider in turn. Note that $O^{p^{\prime}}(\mathcal{F})$ is a simple fusion system supported on $S$.
  • (1)
    $\mathcal{F}$ is not an exotic Benson–Solomon system described in [LO02]: This is immediate, since $p$ is odd.
  • (2)
    $\mathcal{F}$ is not a van Beek exotic fusion system on a subgroup of a Sylow $3$-subgroup of the sporadic Thompson group, or of a Sylow $5$-subgroup of the Monster, as described in [VAN25]: Indeed, $|Z(S)|=q^{2}$ with $q>p$.
  • (3)
    $\mathcal{F}$ is not an exotic fusion system on a $p$-group with an abelian subgroup of index $p$, as described in [OLI14], [COS17], or [OR20]: Since $q>p$, the group $S$ has no such subgroup.
  • (4)
    $\mathcal{F}$ is not an exotic fusion system on a $p$-group of maximal class described in [GP25]: As $q>p$ and $|Z(S)|=q^{2}$, the group $S$ is not of maximal class.
  • (5)
    $\mathcal{F}$ is not a polynomial fusion system on $S_{n}(q)$ for $1\leq n\leq p-1$, as described in [CP10], nor one of its pruned variants described in [HLL23]: In contrast to these $p$-groups, we have $|Z(S)|=q^{2}$ and $S$ has an abelian subgroup of index $q$.
  • (6)
    $\mathcal{F}$ is not a $p^{\prime}$-index exotic subsystem of the $p$-fusion system of a finite simple group $H$ of Lie type in characteristic $r\neq p$, as described in [OR21]: Suppose otherwise, and let $T\in\operatorname{Syl}_{p}(H)$. Then $O^{p^{\prime}}(\mathcal{F}_{T}(H))$ is an exotic fusion system as in [OR21, Theorem A(c)]. In particular, by [OR21, Lemma 2.2(a)],
    $\{V,R^{S}\}\subseteq\mathcal{F}^{rc}=O^{p^{\prime}}(\mathcal{F})^{rc}=\mathcal{F}_{T}(H)^{rc}.$
    Replaying the final two paragraphs of the proof of Proposition 3.8, we conclude that $T\not\cong S$. Hence $O^{p^{\prime}}(\mathcal{F})\not\cong O^{p^{\prime}}(\mathcal{F}_{T}(H))$, as required.
Thus $\mathcal{F}$ is not isomorphic to any of the fusion systems listed above, and these cases exhaust all currently known exotic fusion systems. We conclude that the fusion systems in Theorem 3.9 are new.